OpenStudy (anonymous):

fourth root of 128n to the 8th power

OpenStudy (anonymous):

\[(128n)^2\]

OpenStudy (anonymous):

\[^{} \sqrt[4]{128}n ^{8}\] thats the anwser for this question?

OpenStudy (anonymous):

um. Well, your wording in the original question is rather ambiguous.

OpenStudy (anonymous):

sorry em the question was to simplify \[\sqrt[4]{128}n ^{8}\]

OpenStudy (anonymous):

Ahh

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

\[2*2^{3/4}*n^8=2\sqrt[4]{8}n^8\]

OpenStudy (anonymous):

Or, is the \[n^8\] inside the square root or did you really mean \[\sqrt[4]{128}n^8\]

OpenStudy (anonymous):

if it was what would the answer b i left the paper

OpenStudy (anonymous):

i think it was inside

OpenStudy (anonymous):

ok. Then the answer is \[n^22\sqrt[4]{8}\]

OpenStudy (anonymous):

how many more questions would u be willing to do?